# TRIM: Token Reduction and Inference Modeling for Cost-Effective Language Generation

### v3

@@ Appendix A Supplementary Material
A.1 Exploratory experimentation on the LLM capabilities of ommiting semantic-irrelevant terms
For this exploratory experiment we have used the model "GPT-4o" OpenAI through its API in its version "gpt-4o-2024-08-06" OpenAI (2024). The list of articles and conjunctions used are: the, a, an, as, though, because, before, even, if, that, since, so, than, unless, until, when, whenever, where, whereas, wherever, while, and, but. The set of prompts used in this experimentation are the following:
Prompt 1: Respond just in a paragraph. $\{knowledge\_question\}$
Prompt 2: Respond just in a paragraph in a distilled way removing the redundant and semantic-irrelevant terms of the language such as: $\{list\_of\_terms\}$. That is to say, omit the terms of the ideal answer that do not affect directly to the semantic meaning and can be reconstructed later on based on the context. This task will be carried out by another language model in charge of generating those omitted terms based on the list provided and reconstructing the ideal answer. Return only the distilled answer, nothing else. $\{knowledge\_question\}$
A.2 DKA Dataset elaboration
We have used the model "GPT-4o" OpenAI through its API in its version "gpt-4o-2024-08-06" OpenAI (2024). The 30 general knowledge topics, 20 concepts of each of them and 5 keywords relevant to each concept have been generated using the same LLM with a human curation after generation to verify its validity. Next we show an example of the template questions generation of one keyword (decentralization) associated to one of the concepts (blockchain) related to one of the topics (technology).
@@ A.2 DKA Dataset elaboration
A.3 Consistency among responses in LLMs
Mean Variance
BLEU 0.77 0.03
ROUGE-1 0.87 0.01
ROUGE-L 0.83 0.02
METEOR 0.83 0.02
Cosine Similarity 1.00 0.00
A.4 Aligning Function Words and $\Theta$-metrics
Original The of the of
Reconstructed of the a of
Result omission match insertion substitution match
Table 7: Example of aligned function words and matching result.

### v5

@@ Appendix A Supplementary Material
A.1 Capabilities of LLMS on Omitting Function Words
For this exploratory experiment we have used the model "GPT-4o" OpenAI through its API in its version "gpt-4o-2024-08-06". The list of articles and conjunctions used are: the, a, an, as, though, because, before, even, if, that, since, so, than, unless, until, when, whenever, where, whereas, wherever, while, and, but. The set of prompts used in this experimentation are the following:
Prompt 1: Respond just in a paragraph. $\{query\}$
Prompt 2: Respond just in a paragraph in a distilled way removing the redundant and semantic-irrelevant words of the language such as: $\{list\_of\_words\}$. That is to say, omit the words of the ideal answer that do not affect directly to the semantic meaning and can be reconstructed later on based on the context. This task will be carried out by another language model in charge of generating those omitted words based on the list provided and reconstructing the ideal answer. Return only the distilled answer, nothing else. $\{query\}$
Figure 4 shows the relationship between the length (measured in tokens) of the vanilla answer (answer to prompt 1) and the difference in tokens between the vanilla answer and distilled answer (answer to prompt 2), showing how the trend line grows linearly. It is important to note that both answers were generated in separate sessions, so their differences cannot be fully explained simply by omitted tokens.
Figure 4: Vanilla Answer Length vs Length Difference Between Answers
A.2 NaLDA Dataset Elaboration
To build this dataset, we used two approaches depending on the task type. For knowledge question-answering (QA) and create tasks, we generated synthetic questions using various templates for diversity. For knowledge QA tasks, we created 9,000 questions across 30 general topics (such as Science, History, Art). For each topic, we selected 20 concepts and 5 keywords per concept. We then used templates to generate questions focused on explanations and comparisons between keywords and concepts. The topics, concepts, and keywords were generated with GPT-4o and reviewed by humans for accuracy. Below is an example of a template question using the keyword "decentralization" linked to the concept "blockchain", related to one of the topics "technology".
@@ A.2 NaLDA Dataset Elaboration
For the create tasks, we have gathered 8,800 different instructions related to create a story. Each story creation task is composed of a place, a theme and a historical character from a sample of 20 places, 20 themes and 20 historical characters. These sets have been also generated by the LLM with a human curation after generation to verify its validity. Here is a set of examples of tasks generated:
  • Write a story located in a Pirate Ship, with Abraham Lincoln as the main character and a theme of eroism and Sacrifice.
  • Imagine a tale set in a Industrial Revolution Factory, featuring Napoleon Bonaparte and exploring the theme of Survival and Resilience.
  • Create a narrative in a Underwater City, where Leonardo da Vinci faces challenges related to Family and Legacy.
  • Develop a plot in Santiago Bernabeu, starring Albert Einstein and centered around Fantasy and Magic.
  • Craft a story in a Arctic Base, with Cristiano Ronaldo navigating the complexities of Mystery and Intrigue.
For the re-write, summarize, and translate tasks, we sampled paragraphs from the Wikipedia text corpus and created instructions for each task type, resulting in 9,000 entries per type. For re-write tasks, the LLM was told to change the style. For summarize tasks, we used longer paragraphs to better fit the task and instructed the LLM to summarize them. For translate tasks, we selected Spanish Wikipedia paragraphs and asked the LLM to translate them into English. In our pipeline, each of these entries corresponds to query $Q$.
Furthermore, we have to amend each query with the custom prompt $P^{\prime}$ that instructs the generator model $M_{G}$ to generate both the ideal and distilled version of the answer. This custom prompt $P^{\prime}$ is a variation of the instruction prompt $P$ of the TRIM pipeline where, in addition to generate the distilled answer based on the given set of words $W$, it has to generate also the ideal answer so we can evaluate later on. Ideally, the generation of the ideal and distilled answer should be requested in different sessions, so the distilled one would simulate the real scenario of our pipeline, and the ideal would be the ground truth. Otherwise, the LLM would have the context of the ideal to generate the distilled version. However, after carrying some experiments we realized that doing it in different sessions produce additional variability (e.g., synonyms, grammatical alternative and paraphrasing) that affects directly the evaluation purpose of this dataset, preventing us of using them in the evaluation (see Supplementary Material A.3). Under the assumption that the LLM can generate distilled text directly as shown in Section 3, we keep the single session generation to focus the evaluation on the ability of reconstructing the distilled narrative.
Here is the custom prompt $P^{\prime}$ that is used together with each task to form the dataset: I want you to return only a json with two fields “original_text” and “distilled_text”. The original text is the response to the question in one paragraph. Then, for the distilled text, return the same response but removing the redundant terms of the language such as: $\{\text{set\_of\_easily-inferable\_words}\}$ that do not affect directly to the semantic and can be reconstructed later on based on the context. That is to say, omit the words of the ideal answer that do not affect directly to the semantic meaning and can be reconstructed later on based on the context. This task will be carried out by another language model in charge of generating those omitted words based on the list provided and reconstructing the ideal answer. Keep the rest of the terms and the structure of the original response, only remove those redundant terms. Here is the task: $\{\text{query}\}$
A.3 Consistency Among Responses in LLMs
Mean Variance
BLEU 0.77 0.03
ROUGE-1 0.87 0.01
ROUGE-L 0.83 0.02
METEOR 0.83 0.02
Cosine Similarity 1.00 0.00
A.4 Aligning Function Words and $\Theta$-metrics
Original The of the of
Predict of the a of
Result omission match insertion substitution match
Table 6: Example of aligned function words and matching result.
@@ A.4 Aligning Function Words and $\Theta$-metrics
A.5 Results on Validation Set
Looking at the differences between evaluation sets (NaLDA dataset using GPT-4o as $M_{G}$ and Wikipedia sample), we observe that the results of the $\Theta$ metrics are better on the NaLDA dataset, while the rest of the metrics are slightly better on average in the Wikipedia sample (See Table 8). This indicates that, while the reconstruction model generally performs better on the validation dataset—built with the same philosophy as the training data—it achieves better results on metrics focused solely on reconstructing the pre-defined word set when the LLM chooses which terms to omit during generation, rather than omitting all of them directly.
Furthermore, an interesting insight emerges in the $\Theta$ metrics of the distilled text (i.e., matching distilled text with the ground truth). In the Wikipedia sample evaluation, the $\Theta$-recall is 0 because these words are removed directly from the text. In contrast, in the NaLDA dataset evaluation, the $\Theta$-recall is $8.6\%$, showing that the LLM does not erase all words from the list, but only those it considers easily-inferable from context, as instructed in the prompt (see Appendix A.2).
Both of these insights demonstrate a desirable capability of the LLM: not only omitting a pre-defined set of words during generation, but also identifying when a word is not easily-inferable based on context and choosing not to omit it, even if it is in the pre-defined word set $W$. These results are consistent with the exploratory analysis shown in Table 1.
@@ A.5 Results on Validation Set
Figure 5: Visual comparison of the results for the reconstruction models in the NaLDA dataset using GPT-4o as $M_{G}$
Metric Bart T5 Qwen2.5-0.5 Qwen2.5-1.5
GPT-4o
$\Theta$ Precision (%) 88.81(9.59) 90.14(9.41) 82.11(11.64) 85.35(10.78)
$\Theta$ Recall (%) 82.50(11.04) 82.24(11.91) 84.33(10.27) 87.65(9.19)
$\Theta$ F1 (%) 85.11(9.37) 85.51(9.79) 82.77(9.98) 86.10(9.12)
SacreBLEU (%) 81.32(9.83) 82.92(10.69) 81.59(9.84) 83.99(9.25)
METEOR (%) 91.51(5.79) 92.15(6.85) 92.73(5.42) 93.70(4.84)
ROUGE-1 (%) 94.82(3.33) 95.13(3.99) 94.62(3.54) 95.26(3.25)
ROUGE-L (%) 93.56(4.08) 93.96(4.66) 93.03(4.43) 94.10(3.98)
Cosine Similarity (%) 99.98(0.03) 99.99(0.02) 99.98(0.02) 99.98(0.03)
Perplexity 35.06(18.72) 31.53(17.46) 32.52(17.48) 31.13(16.54)
Perplexity Original 25.59(13.87) 25.59(13.87) 25.58(13.84) 25.58(13.85)
Saved Tokens (%) 19.41(5.84) 19.41(5.84) 19.41(5.83) 19.41(5.83)
GPT-4.1-mini
$\Theta$ Precision (%) 86.29(11.35) 87.14(11.28) 78.61(13.06) 82.13(12.27)
$\Theta$ Recall (%) 79.79(11.95) 79.64(12.39) 81.98(11.23) 85.60(10.29)
$\Theta$ F1 (%) 82.28(10.60) 82.52(10.87) 79.64(11.08) 83.25(10.31)
SacreBLEU (%) 73.16(10.02) 75.21(10.33) 73.29(10.01) 76.02(9.64)
METEOR (%) 87.59(5.73) 88.49(6.01) 88.68(5.50) 89.85(5.13)
ROUGE-1 (%) 92.07(3.42) 92.58(3.68) 91.60(3.65) 92.31(3.50)
ROUGE-L (%) 90.30(4.16) 90.95(4.35) 89.45(4.51) 90.72(4.20)
Cosine Similarity (%) 99.98(0.03) 99.98(0.02) 99.97(0.03) 99.97(0.03)
Perplexity 42.06(21.90) 37.67(20.43) 39.30(20.42) 37.05(19.42)
Perplexity Original 26.52(16.33) 26.51(16.13) 26.40(15.46) 26.35(14.81)
Saved Tokens (%) 22.80(5.44) 22.79(5.44) 22.80(5.43) 22.81(5.42)
GPT-4.1
$\Theta$ Precision (%) 87.73(10.19) 88.41(10.10) 79.53(12.38) 83.20(11.51)
$\Theta$ Recall (%) 82.31(11.65) 82.22(12.01) 84.09(11.00) 87.51(9.91)
$\Theta$ F1 (%) 84.46(9.79) 84.73(10.06) 81.23(10.56) 84.84(9.65)
SacreBLEU (%) 79.66(9.57) 81.78(9.48) 79.23(9.60) 81.60(9.17)
METEOR (%) 91.11(5.14) 92.10(5.00) 92.07(4.89) 92.93(4.58)
ROUGE-1 (%) 94.30(3.15) 94.86(3.09) 93.79(3.39) 94.42(3.25)
ROUGE-L (%) 93.09(3.80) 93.75(3.70) 92.28(4.16) 93.31(3.88)
Cosine Similarity (%) 99.98(0.02) 99.99(0.02) 99.98(0.02) 99.98(0.03)
Perplexity 38.84(19.61) 35.03(18.20) 36.43(18.41) 34.96(17.52)
Perplexity Original 27.18(12.89) 27.18(12.89) 27.18(12.89) 27.18(12.89)
Saved Tokens (%) 19.45(5.32) 19.45(5.32) 19.45(5.32) 19.46(5.32)
Sonnet-4
$\Theta$ Precision (%) 86.46(12.59) 87.15(12.42) 77.68(14.17) 81.53(13.39)
$\Theta$ Recall (%) 82.75(12.97) 81.56(13.74) 84.01(12.53) 87.82(11.44)
$\Theta$ F1 (%) 83.73(12.09) 83.39(12.51) 79.79(12.66) 83.70(11.88)
SacreBLEU (%) 84.36(8.43) 85.13(9.00) 83.56(8.53) 86.25(7.84)
METEOR (%) 93.61(4.36) 93.48(5.26) 94.48(4.15) 95.43(3.55)
ROUGE-1 (%) 95.73(2.72) 95.71(3.28) 95.33(2.95) 96.01(2.67)
ROUGE-L (%) 94.81(3.23) 94.81(3.74) 93.95(3.64) 95.11(3.16)
Cosine Similarity (%) 99.99(0.03) 99.99(0.02) 99.98(0.02) 99.98(0.02)
Perplexity 38.04(19.73) 35.36(19.04) 35.86(18.54) 36.52(223.85)
Perplexity Original 30.84(18.05) 30.85(18.06) 30.79(17.60) 30.81(17.83)
Saved Tokens (%) 17.83(5.00) 17.83(5.00) 17.83(4.99) 17.83(4.99)
Opus-4.1
$\Theta$ Precision (%) 88.17(10.03) 88.56(9.98) 79.26(12.32) 83.34(11.37)
$\Theta$ Recall (%) 83.19(11.23) 81.12(12.12) 84.53(10.76) 88.56(9.41)
$\Theta$ F1 (%) 85.05(9.75) 84.03(10.32) 81.15(10.73) 85.27(9.71)
SacreBLEU (%) 86.56(7.35) 86.49(8.25) 85.85(7.43) 88.61(6.69)
METEOR (%) 94.51(4.01) 93.76(5.17) 95.59(3.68) 96.54(3.01)
ROUGE-1 (%) 96.50(2.40) 96.11(3.09) 96.26(2.60) 96.94(2.24)
ROUGE-L (%) 95.62(2.86) 95.22(3.50) 94.89(3.25) 96.06(2.73)
Cosine Similarity (%) 99.98(0.03) 99.99(0.02) 99.98(0.02) 99.98(0.02)
Perplexity 37.48(18.85) 35.25(18.36) 35.28(17.95) 33.78(17.07)
Perplexity Original 30.74(15.90) 30.74(15.90) 30.75(15.90) 30.75(15.89)
Saved Tokens (%) 17.38(4.36) 17.38(4.36) 17.38(4.36) 17.38(4.37)
Table 9: Average (standard deviation) metrics results over the NaLDA dataset for each $M_{G}$ and $MR$ models using the word set $W$ obtained during evaluation.